find the length of a chord of parabola passing through the vertex and making an angle theta with the x-axis

Grade:

1 Answers

Aman Bansal
11 years ago

Dear Sai,

Let F(h, k) be the focus and the equation ax + by + c = 0 be the equation of the directrix of a parabola. Let P(x, y) be any point on the parabola and let PM be the perpendicular drawn from p on the directrix .

Then PF = PM => √(x - h)2 + (y - k)= ax + by + c/√a+ b2

=> (x - h)2 + (y - k)= (ax + by + c)2/a2 + b2

=>a2 + b2)[(x - h)+ (y - k)2] = (ax + by + c)2

=> (bx - ay)2 + 2(-ha- hb2 - ac) x + 2( - ka- kb- bc)y + (h2a+ h2b+ k2a+ k2b- c2) = 0

This is of the form (bx - ay)+ 2gx + 2fy + d = 0.

Thus , the general form of parabola equation is a second degree equation in which second degree terms form a perfect square. Solve parabola equation in the above method. Also for more help one can connect to an online tutor and get the required help.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert